Presented below are selected transactions on the books of Cullumber Corporation.

May 1, 2017 Bonds payable with a par value of $861,600, which are dated January 1, 2017, are sold at 106 plus accrued interest. They are coupon bonds, bear interest at 11% (payable annually at January 1), and mature January 1, 2027. (Use interest expense account for accrued interest.)
Dec. 31 Adjusting entries are made to record the accrued interest on the bonds, and the amortization of the proper amount of premium. (Use straight-line amortization.)
Jan. 1, 2018 Interest on the bonds is paid.
April 1 Bonds with par value of $344,640 are called at 102 plus accrued interest, and redeemed. (Bond premium is to be amortized only at the end of each year.)
Dec. 31 Adjusting entries are made to record the accrued interest on the bonds, and the proper amount of premium amortized.


Prepare journal entries for the transactions above. (Round intermediate calculations to 6 decimal places, e.g. 1.251247 and final answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

May 1, 2017Dec. 31, 2017Jan. 1, 2018April 1, 2018Dec. 31, 2018

May 1, 2017Dec. 31, 2017Jan. 1, 2018April 1, 2018Dec. 31, 2018

(To record the interest)

(To amortize the premium)

May 1, 2017Dec. 31, 2017Jan. 1, 2018April 1, 2018Dec. 31, 2018

Apr. 1, 2018

May 1, 2017Dec. 31, 2017Jan. 1, 2018April 1, 2018Dec. 31, 2018

(To record the interest)

(To amortize the premium)
Answer #1
Date Account title Debit credit
May 1 2017 cash 944888
Interest expense 31592
Premium on bond payable 51696
Bond payable 861600
Dec 31 Interest expense (861600*11%) 94776
Interest payable 94776
Dec 31 Premium on bond payable 3565
Interest expense 3565
Jan 1 2018 Interest payable 94776
cash 94776
April 1 Bond payable 344640
Premium on bond payable 18175
Interest expense 9478
Cash (344640*102/100) 351533
Gain on redemption of bond 20760
Dec 31 Interest expense 57966
Interest payable (526960*11%) 57966
Dec 31 Premium on bond payable 3795
Interest expense (3276+519 ) 3795

working:

May 1)Issue price = 861600*106 /100 = 913296

Premium on bond payable = 913296 -861600 = 51696

Interest is accrued for 4 months (1Jan -1May ) = 861600*11%*4/12 = 31592

Dec 31)Years to maturity = 1 May 2017 - 1 Jan 2027 = 116 months

Amortization of premium per month = 51696 /116 =$ 445.655172 per month

Amortization on Dec 31 2017 = 8 months *445.65517 = 3565 [1 May -31 Dec]

1 April 2018 )

Interest is accrued for 3 months : 344640*11%*3/12 = 9478

Un-amortized Bond premium as on 1 Jan 2018

51696 - 3565

48131 on 861600 par value bonds

Total amortization for 2018

Premium on 334640 bond to retired

48131*334640/861600

18694

premium on remaining 861600-334640 retired= 526960 bonds

48131 *526960/861600

29437

3 months amortization on bond retired: 18694 *3/108 = 519 per month amortization of bond premium :29437 /108 = 273 or 273*12= 3276 3795
Premium remaining : 18694 -519= 18175

Months left to maturity (1 Jan 2018 -1J an 2027 ) = 9 years *12 = 108